A golf ball is hit upward with a velocity of 66 feet per second (45 mph), its height h in feet above the ground after t seconds can be modeled by .
a. Determine when the ball strikes the ground.
b. When is the height of the ball 50 feet?




Answer by stanbon(75887)

A golf ball is hit upward with a velocity of 66 feet per second (45 mph), its height h in feet above the ground after t seconds can be modeled by
h(t) = -16t^2+66t
----------------------------.
a. Determine when the ball strikes the ground.
The height will be zero when the ball trikes the ground:
Solve:
-16t^2+66t = 0
---
Factor:
(-2t)(8t-33) = 0
t = 0 before the ball is hit.
t = 33/2 = 15.5 seconds (time it hits the ground)
------

b. When is the height of the ball 50 feet?
Solve:
-16t*2 + 66t = 50 feet
-16t^2 + 66t - 50 = 0
-2(8t^2 - 32t + 25) = 0
-----
8t^2-32t+25 = 0
------
t = [32 +- sqrt(32^2-4*8*25)]/16
------
t = [32 +- sqrt(224)]/16
---
t = [32 +- sqrt(4*56)]/16
----
t = 1.021 sec. on the way up
t = 1.065 sec. on the way down
=================
Cheers,
Stan H.
=================